Consider the pdf of a Normal variable 𝑿 is defined as 𝒇(𝒙) = (1/√98𝝅)e^-((x+11)^2/98))

i) Find the mgf of 𝑿.

ii) Evaluate 𝑷(βˆ’πŸ‘ > βˆ’π‘Ώ > πŸπŸ‘).

iii) Find the value of π‘ͺ such that 𝑷(|𝑿 + 𝟏𝟏| β‰₯ π‘ͺ) = 𝟎. πŸŽπŸ”πŸπŸ’.

iv) Find βˆ’π’πŸŽ.πŸŽπŸŽπŸ‘πŸ and convert to 𝑿.

1
Expert's answer
2021-10-03T17:44:52-0400

i)

The moment generating function (MGF):

"m_X(t)=Ee^{tX}"

"m_X(t)=\int^{\infin}_{-\infin}e^{tx}f(x)dx=\int^{\infin}_{-\infin}\frac{e^{tx}e^{-(x+11)^2/98}}{\sqrt{98\pi}}dx="


"=\frac{e^{-121/98}}{\sqrt{98\pi}}\int^{\infin}_{-\infin}e^{-(x^2+x(98t+22))/98}dx=\frac{e^{-121/98}}{\sqrt{98\pi}}\cdot e^{(98t+22)^2/392}\sqrt{98\pi}=e^{(24.5t^2+11t)}"


ii)

"P(-3>-X>13)=\int^{-3}_{-\infin}f(-x)dx+\int^{\infin}_{13}f(-x)dx="


"=\int^{-3}_{-\infin}\frac{e^{-(-x+11)^2/98}}{\sqrt{98\pi}}dx+\int^{\infin}_{13}\frac{e^{-(-x+11)^2/98}}{\sqrt{98\pi}}dx=-\frac{(erf(14/\sqrt{98})-1)\sqrt{98\pi}}{2\sqrt{98\pi}}-\frac{(erf(2/\sqrt{98})-1)\sqrt{98\pi}}{2\sqrt{98\pi}}"

where erf is error function.

"erf(14/\sqrt{98})=0.95,\ erf(2/\sqrt{98})=0.22"

"P(-3>-X>13)=0.025+0.390=0.415"


iii)

"P(|X+11|\ge C)=2\int^{\infin}_{C-11}\frac{e^{-(x+11)^2/98}}{\sqrt{98\pi}}dx=0.0614"


"2\cdot-\frac{erf((\sqrt{2}(C-11)+11\sqrt{2})/14)-1}{2}=0.0614"

"erf((\sqrt{2}(C-11)+11\sqrt{2})/14)=1-0.0614=0.9386"

"(\sqrt{2}(C-11)+11\sqrt{2})/14=1.323"


"C=\frac{14\cdot 1.323-11\sqrt{2}}{\sqrt{2}}+11=13.1"


iv)

"Z_{0.0031}=-3.42"

"-Z_{0.0031}=3.42"

"\frac{x-\mu}{\sigma}=3.42"

"\mu=-11,\ \sigma=7"

"X=3.42\sigma+\mu=3.42\cdot7-11=12.94"